Find the indefinite integrals for the following:

a.) \int \frac{2}{9x+1}dx

b.) \int \frac{2x}{9x^2+1}dx

c.) \int (lnx)^2\cdot \frac{1}{x}dx

d.) \int xe^9^x^{2}^+^1dx (the exponent is 9x^2+1 )

Please show work, thanks!
Answer #1

9 T- JAEG 47 += 9x+1 3 dx = 学 5) Ia en 141 +0 |-j Angx+1 + | (t = 9xt+1 紫=184 dy = 18 n da : T = 了全一是 日 12 | take the 中 |-- j

(9 xv + z(xur) -I let y= enx 7 dx= x.dy fp** CM) S =1 hp zh s = I E = I= /2 y + c » 1= {(4x32 +] let y= 9x²+1 * dx = da :: 1=